Conduct a horizontal analysis (comparative analysis) for the balance sheets of Gian’s Italian Cuisine as of...

  1. Conduct a horizontal analysis (comparative analysis) for the balance sheets of Gian’s Italian Cuisine as of November 30 and December 31, 2016.

November

December

Current Assets:

Cash

9,000

14,000

Accounts Receivable

10,000

11,500

Food Inventory

4,000

3,538

Beverage Inventory

6,000

3,950

Prepaid Rent

18,000

20,000

     Total Current Assets

Fixed Assets:

Furniture, Fixture & Equipment (FF&E)

7,500

9,870

Accumulated Depreciation (FF&E)

(3,030)

(4,000)

Building

20,000

20,000

Long term Investments

5,000

8,580

     Long Term Assets

29,470

34,450

Total Assets

Current Liabilities:

Accounts Payable

7,000

6,790

Accrued Payroll

5,400

6,540

Accrued Taxes

5,500

3,500

     Total Current Liabilities

Long Term Debt

10,000

12,000

     Total Liabilities

Owner's Equity

39,000

40,620

Retained Earnings

9,570

17,988

    Total Owner's Equity

Total Liabilities and Owner's Equity
Answer #1

Horizontal Analysis (Comparative Analysis) for the balance sheet of Gian's Italian Cuisine as of November 30 & December 31, 2016 :

November December Increase or (Decrease)
Amount Percent
Current Assets:
Cash 9,000 14,000 5,000 55.56% = (5,000/9,000) * 100
Accounts Receivable 10,000 11,500 1,500 15% = (1,500/10,000) * 100
Food Inventory 4,000 3,538 -462 (11.55%) = (-462/4,000) * 100
Beverage Inventory 6,000 3,950 -2,050 (34.17%) = (-2,050/6000) * 100
Prepaid Rent 18,000 20,000 2,000 11.11% = (2,000/18,000) *100
Total Current Assets 47,000 52,988 5,988 12.74% = (5,988/47,000) * 100
Fixed Assets:
Furniture, Fixture & Equipment (FF&E) 7,500 9,870 2,370 31.6% = (2,370/7,500) * 100
Accumulated Depreciation (FF&E) -3,030 -4,000 -970 32.01% = (-970 / -3,030) * 100
Building 20,000 20,000 0 0%
Long Term Investments 5,000 8,580 3,580 71.6% = (3,580/5,000) * 100
Long Term Assets 29,470 34,450 4,980 16.90% = (4,980/29,470) * 100
Total Assets 76,470 87,438 10,968 14.34% = (10,968/79,470) * 100
Current Liabilities:
Accounts Payable 7,000 6,790 -210

(3%) = (-210/7,000) * 100

Accrued Payroll 5,400 6,540 1,140 21.11% = (1,140/5,400) * 100
Accrued Taxes 5,500 3,500 -2,000 (36.36%) = (-2,000/5,500) * 100
Total Current Liabilities 17,900 16,830 -1,070 (5.98%) = (-1,070/17,900) * 100
Long Term Debt 10,000 12,000 2,000 20% = (2,000/10,000) * 100
Total Liabilities 27,900 28,830 930 3.33% = (930/27,900) * 100
Owner's Equity 39,000 40,620 1,620 4.15% = (1,620/39,000) * 100
Retained Earnings 9,570 17,988 8,418 87.96% = (8,418/9,570) * 100
Total Owner's Equity 48,570 58,608 10,038 20.67% = (10,038/48,570) * 100
Total Liabilities and Owner's Equity 76,470 87,438 10,968 14.34% = (10,968/76,470) *100

Percentage Change in Increase / (Decrease) = [Amount of Increase or (Decrease) / Amount on November 30, 2016] * 100

Percentage Change in Increase / (Decrease) is Rounded off to 2 Decimals.